Hello my name is ChristopherI am hoping you can help me on this HW problem...

Solve equation by putting in quadratic form (Use a U substitution)....?
Very confused could you please help me ? thankyou
4x^4-7x^2+3=0

Answer by stanbon(75887)

Solve equation by putting in quadratic form (Use a U substitution)....?
Very confused could you please help me ? thankyou
4x^4-7x^2+3=0
----
Let u = x^2
---
Substitute to get:
4u^2-7u+3 = 0
----
Factor:
4u^2 -4u-3u + 3 = 0
---
4u(u-1)-3(u-1) = 0
---
Factor Again:
(u-1)(4u-3) = 0
---
u = 1 or u = 4/3
-----
Convert answer to "x":
x^2 = 1 or x^2 = 4/3
---
x = +-1 or x = +-2/sqrt(3)
==============================
Cheers,
Stan H.